# Shen2018S18

### v1

@@ Abstract
A $33\,M_{\odot}$ black hole (BH) was recently discovered in an 11.6-year binary only 590 pc from the Sun. The system, Gaia BH3, contains a $0.8\,M_{\odot}$ low-metallicity giant ($\rm[M/H]=-2.2$) and is kinematically part of the Galactic halo, suggesting that the BH formed from a low-metallicity massive star. I show that orbits similar to that of Gaia BH3 are naturally produced through isolated binary evolution. The system’s period and eccentricity can result from a broad range of initial orbits with a modest natal kick ($v_{\rm kick}\lesssim 75\,\rm km\,s^{-1}$) to the BH. I construct MESA models for metal-poor massive stars with initial masses ranging from $35-55\,M_{\odot}$, which reach maximum radii of $1150-1800\,R_{\odot}$ as red supergiants. Stars of this size would fit inside most plausible pre-supernova orbits for the system without overflowing their Roche lobes. In addition, models with moderately rapid initial rotation ($\Omega/\Omega_{\rm crit}\gtrsim 0.45$) undergo chemically homogeneous evolution and never expand to radii larger than $10\,R_{\odot}$. There are thus multiple channels through which a low-metallicity, extreme-mass ratio binary could produce a system like Gaia BH3. Dynamical formation scenarios are also viable, and there is little doubt that both isolated and dynamically-formed BH binaries with orbits similar to Gaia BH3 will be discovered in Gaia DR4. Only about 1 in 10,000 stars in the solar neighborhood have metallicities as low as Gaia BH3. This suggests that BH companions are dramatically over-represented at low-metallicity, though caveats related to small number statistics apply. The fact that the luminous star in Gaia BH3 has been a giant — greatly boosting its detectability — only for $\sim$1of the time since the system’s formation implies that additional massive BHs remain to be discovered with only moderately fainter companions.
@@ 1 Introduction
Gaia Collaboration et al. (2024) recently reported discovery of an astrometric binary containing a $0.8\,M_{\odot}$ star on the lower giant branch and a $33\,M_{\odot}$ dark companion that is presumably a black hole (BH). The system, which they named Gaia BH3, was discovered with pre-release DR4 astrometry, and its orbital motion was confirmed with (nearly) independent radial velocity measurements from the Gaia RVS spectrometer.
@@
Gaia BH3 is the widest BH binary discovered to date, with an orbital period of $\sim 4250$ days and a semimajor axis of $\sim 16.5$ au. The eccentricity is relatively high, $e\approx 0.73$, such that the star and BH pass within 4.5 au of each other at periastron. How the system formed is uncertain. Gaia Collaboration et al. (2024) note that a typical red supergiant would not fit within the current orbit at periastron. They thus consider formation of the system from an isolated binary unlikely, and invoke dynamical exchange as a possible alternative.
Here, I consider the possibility of formation from an isolated binary in more detail. I conclude that – although dynamical formation channels are also viable – Gaia BH3’s properties may be naturally explained as the end state of an isolated, extreme mass ratio binary. The main differences between my analysis and that of Gaia Collaboration et al. (2024) are that (a) I consider the possibility that the orbit today may be different from the orbit at the time of the BH’s formation, due to kicks and/or mass loss during the BH progenitor’s death, and (b) I show that low-metallicity massive stars remain compact for most of their evolution, expanding to red supergiant dimensions only in the final $\sim 1000$ years of their lives, if at all. As a result, there is a broad parameter space of initial orbits in which the BH progenitor either never fills its Roche lobe, or fills it too late for the companion star to spiral inward significantly.
@@
The remainder of this paper is organized as follows. Section 2 uses Monte Carlo simulations of BH formation to assess the likely initial separation if Gaia BH3 formed from an isolated binary. In Section 3, I describe evolutionary models for low-metallicity massive stars that could form the BH. Section 4 argues that BH companions appear to be strongly overrepresented at low metallicity, and that additional nearby BHs orbited by low-metallicity stars likely remain to be discovered. Future prospects and broader implications are briefly discussed in Section 5.
2 Constraints on the progenitor orbit
I use Monte Carlo simulations to study the combinations of pre-supernova [^1] Here “supernova” loosely refers to the death of the massive star and formation of the BH. Whether there was a successful explosion or a long-lived luminous transient is uncertain. (SN) orbits and natal kicks that could have produced an orbit similar to that of Gaia BH3. This approach follows Tauris et al. (2017) and El-Badry et al. (2024); see Tauris & van den Heuvel (2023) for additional details about SN kicks. In brief, I simulate a large number ($N=10^{7}$) of orbits and kicks, and then use the formalism from Brandt & Podsiadlowski (1995) to predict the post-SN parameters. Finally, I analyze the initial parameters of the simulations for which the final period and eccentricity are close to the observed values for Gaia BH3: $P_{\rm orb,final}=4000-4500$ d, and $e_{\rm final}=0.68-0.78$. Narrowing these ranges reduces the number of surviving Monte Carlo samples but has no significant effect on their distribution.
Figure 1: Constraints on kicks and the pre-SN orbit of Gaia BH3-like binaries. I assume that the binary had a circular orbit before the SN and that the BH received a kick with velocity $v_{\rm kick}$ when it formed. I simulate $10^{7}$ pre-SN orbits with a uniform period distribution $P_{\rm orb}\sim\mathcal{U}(0,2P_{\rm orb,BH3})$, a pre-SN mass distribution $M_{2,\rm\,init}/M_{\odot}\sim\mathcal{U}(33,53)$, and a kick velocity distribution $v_{\rm kick}/({\rm km\,s^{-1}})\sim\mathcal{U}(0,100)$. I then show properties of the systems whose post-SN orbits have periods and eccentricities similar to Gaia BH3 (upper left). These orbits have a wide range of pre-SN separations, ranging from 1000 to 6000 $R_{\odot}$.
I begin with a uniform distribution of pre-SN orbital periods between 0 and 8500 d (twice the currently observed orbital period of Gaia BH3). I simulate kicks assuming orientations distributed uniformly on the sky and velocities distributed uniformly between 0 and 100 $\rm km\,s^{-1}$. I take a uniform distribution of $\Delta m$, the mass loss of the BH progenitor during the SN, between 0 and 20 $M_{\odot}$, implying a progenitor mass of 33-53 $M_{\odot}$. I assume the pre-SN orbit had $e=0$, having been circularized by tides when the BH progenitor expanded. This is not a population synthesis simulation – pre-SN orbits are unlikely to be uniformly distributed in all parameters – but should be viewed as an exploration of what combinations of orbits, kicks, and mass loss could have produced the observed orbit. Later, I will check whether the possible progenitor orbits are consistent with the expected size of the BH progenitor as a red supergiant. If not – or if an orbit similar to the observed one could only be produced self-consistently from very fine-tuned initial conditions – this would suggest that formation from an isolated binary is unlikely.
The results of this experiment are shown in Figure 1. Initial separations ranging from $\sim 800$ to $\sim 6000\,R_{\odot}$ can produce the observed orbit. Initial separation closer than $\sim 800\,R_{\odot}$ cannot, because widening these to the separation observed today via a kick or mass loss would result in a higher final eccentricity than is observed. Wider initial separations could match the orbit for suitably chosen kicks, but these are excluded by our prior of $P_{\rm orb,\,init}<8500$ d. The median and middle 68% range of pre-SN separations that reproduce the observed orbit is $a_{\rm init}=4473_{-2146}^{+1347}\,R_{\odot}$; this range encompasses the semimajor axis observed today, which is $\approx 3350\,R_{\odot}$.
A broad range of $v_{\rm kick}$ can also reproduce the observed orbit; I find a middle 68% range of $v_{\rm kick}=39_{-14}^{+19}\,\rm km\,s^{-1}$. There no strong correlation between the allowed $v_{\rm kick}$ and $a_{\rm init}$, but the combination of low $v_{\rm kick}$ ($\lesssim 20\,\rm km\,s^{-1}$) and large initial separation ($\gtrsim 1200\,R_{\odot}$) is ruled out because such orbits produce too low eccentricities. The pre-SN mass is not strongly constrained. Only for $M_{2,\rm init}>67\,M_{\odot}$ (where most orbits become unbound due to Blaauw 1961 kicks) would this parameter strongly affect the results. In the limit of $M_{2,\rm init}=33\,M_{\odot}$ (no mass loss), the required initial separation to explain the orbit decreases slightly, to $a_{{\rm init}}=4010_{-2539}^{+1314}R_{\odot}$.
Assuming the system formed from an isolated binary and the orbit was affected by a natal kick, we can ask whether significant fine-tuning is required to explain the orbit. Fine-tuning is always required to produce any specific orbit; here I am concerned with whether the orbit of Gaia BH3 is typical of the predicted BH + main sequence star population, given the inferred kick velocities. If I fix $v_{\rm kick}=39\,\rm km\,s^{-1}$ and repeat the experiment, I find that $\approx 50\%$ of all simulated orbits remain bound, and the median eccentricity of the orbits that remain bound is 0.68. This suggests that Gaia BH3 is typical of the predicted population and the orbit does not require much fine-tuning to explain. If $v_{\rm kick}$ is increased to $80\,\rm km\,s^{-1}$, the fraction of orbits that remain bound falls to 20%.
The largest star that could fit inside a pre-SN orbit of semimajor axis $a_{\rm init}$ has radius $R_{\rm max}=a_{\rm init}f_{q}$, where $f_{q}(M_{1}/M_{2})$ is given by Eggleton (1983) and varies from 0.68 to 0.70 over $M_{1}=35-50\,M_{\odot}$, assuming $M_{2}=0.8\,M_{\odot}$. This means for the pre-SN orbits that reproduce the observed period and eccentricity, the ±$\pm 1\sigma$ range of maximum radii is $R_{\rm max}\approx 3090_{-2170}^{+930}\,R_{\odot}$. We are now in a position to ask whether the progenitor of the BH, which was presumably a low-metallicity massive star, would have fit inside the pre-SN orbit.
@@ 3 The BH progenitor
Figure 2: MESA evolutionary models for low-metallicty massive stars. The model shown in black is the same in all three panels: a $35\,M_{\odot}$ star with metallicity $\rm[M/H]=-2.2$ and no rotation. Colored lines show models varying metallicity (upper left), initial mass (upper right), initial rotation rate (lower left), and overshooting or mixing length (lower right). Diamonds, circles, and stars mark core hydrogen exhaustion, the onset of carbon burning, and core collapse. Dashed diagonal lines show constant radius. Shaded gray region marks $R>3000\,R_{\odot}$, beyond which the star would have been unlikely to fit inside the pre-SN orbit of Gaia BH3. All of the models avoid this limit, and some remain much smaller than it, such that they would not overflow their Roche lobes for any plausible progenitor orbit of Gaia BH3. Lower left highlights models undergoing chemically homogeneous evolution (CHE), which stay smaller than $10\,R_{\odot}$ at this metallicity, even for modest initial rotation rates, $\Omega/\Omega_{\rm crit}\gtrsim 0.45$. CHE represents one possible channel to form Gaia BH3, but even models without rotation could quite plausibly fit within the pre-SN orbit.
I calculated a small suite of evolutionary models for the progenitor of the BH using MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). The calculations closely follow the setup described by Klencki et al. (2020), with MESA updated to version r23.05.1. The most important ingredients are summarized below, and I refer to Klencki et al. (2020, 2021) for more details.
@@
To verify that the problem setup is consistent with the one used by Klencki et al. (2020), I began by reproducing their $35\,M_{\odot}$ at metallicity $Z=0.00017$ and comparing to their published tracks. The model reaches a maximum radius of $1331\,R_{\odot}$ shortly before core collapse. Their corresponding track has a maximum radius $1340\,R_{\odot}$ at the same point, and its evolution in the HR diagram is nearly indistinguishable from that of the model calculated here.
@@
All the models shown in Figure 2 are consistent with fitting within the observed orbit for more than half of the sampled initial separations: the shaded regions in Figure 2 show $R>3000\,R_{\odot}$, above which stars would overflow their Roche lobes in Gaia BH3 for about half of the possible initial orbits shown in Figure 1. None of the models reach radii of $3000\,R_{\odot}$, implying that they would not overflow their Roche lobes for a majority of the possible pre-SN orbits.
One of the main differences between these and higher-metallicity models is that these models remain compact during core He burning: although they are red supergiants at the end of their lives, they spend the large majority of their post-main sequence evolution as blue or yellow supergiants with $R\lesssim 100\,R_{\odot}$, and only expand to red supergiant dimensions during their final $\sim 1000$ years (see Klencki et al., 2020, for further discussion). This is illustrated in Figure 3, which shows the late-stage radius evolution for models with a range of masses (top panel) and metallicities (bottom panel). The fiducial model increases its radius by a factor of 2 in its final 800 years. This behavior is not found in higher metallicity models, even at “low” metallicites comparable to those found in the Magellanic clouds. The upshot is that if a star with properties similar to those predicted by the MESA models overflowed its Roche lobe in the Gaia BH3 system, it would only have done so in the final few hundred years of its evolution.
Figure 3: Radius evolution of MESA models in the final years before core collapse. Mass is varied in the top panel, and metallicity in the bottom panel. At low metallicity, the models expand significantly in the final $\sim 1000$ yr of their evolution, only appearing as red supergiants shortly before core collapse. Expansion occurs earlier at higher metallicity. If the fiducial model (black) were put in an orbit in which it overflowed its Roche lobe at $R=1000\,R_{\odot}$, this would occur only a few tens of orbits before core collapse, giving the companion insufficient time to spiral inward very far.
If a red supergiant overflows its Roche lobe, the inspiral of the companion star is not instantaneous, but occurs only once the work done by the drag force is comparable to the star’s orbital energy. For the large mass ratios and diffuse envelopes of interest here, the expected inspiral time from the outer envelope is of order $10^{2}$ orbits (O’Connor et al., 2023), or of order $10^{3}$ years. Since most of our models expand significantly in their final $\sim 1000$ years, it is possible that the BH progenitor did overflow its Roche lobe, but then collapsed soon after, not leaving the companion star enough time to spiral in. Such an evolutionary history is most likely to be tenable for pre-SN orbits in which the BH progenitor just barely overflows its Roche lobe.
@@
The lower left panel of Figure 2 shows another possible formation channel: formation from a rapidly-rotating star through chemically homogeneous evolution (CHE). At high rotation rates, large-scale meridional circulations efficiently mix helium into the envelope, preventing the buildup of a chemical gradient and core/envelope dichotomy that leads to expansion in nonrotating models (Maeder, 1987; Yoon & Langer, 2005; Woosley & Heger, 2006). Under CHE, stars evolve from the main sequence toward the helium main sequence, where they have effective temperatures above $10^{5}\,\rm K$ and radii smaller than $1\,R_{\odot}$. These models eventually ignite carbon burning and terminate their evolution without ever reaching $R>10\,R_{\odot}$. At the low metallicities considered here, CHE occurs with modest initial rotation rates, $\Omega/\Omega_{\rm crit}\geq 0.45$.

### v2

@@ Abstract
A $33\,M_{\odot}$ black hole (BH) was recently discovered in an 11.6-year binary only 590 pc from the Sun. The system, Gaia BH3, contains a $0.8\,M_{\odot}$ low-metallicity giant ($\rm[M/H]=-2.2$) that is a member of the ED-2 stellar stream. This paper investigates whether the system could have formed via isolated binary evolution. I construct evolutionary models for metal-poor massive stars with initial masses ranging from $35-55\,M_{\odot}$, which reach maximum radii of $1150-1800\,R_{\odot}$ as red supergiants. I then explore what combinations of initial orbit, mass loss, and natal kick can produce the period and eccentricity of Gaia BH3. Initial orbits wide enough to accommodate the BH progenitor as a red supergiant can match the observed period and eccentricity, but only if the BH formed with a significant natal kick ($v_{\rm kick}\gtrsim 20\,{\rm km\,s^{-1}}$). These models are disfavored because such a kick would have ejected the binary from the ED-2 progenitor cluster. I conclude that Gaia BH3 likely formed through dynamical interactions, unless the BH progenitor did not expand to become a red supergiant. Only about 1 in 10,000 stars in the solar neighborhood have metallicities as low as Gaia BH3. This suggests that BH companions are dramatically over-represented at low-metallicity, though caveats related to small number statistics apply. The fact that the luminous star in Gaia BH3 has been a giant – greatly boosting its detectability – only for $\sim$1of the time since the system’s formation implies that additional massive BHs remain to be discovered with only moderately fainter companions. Both isolated and dynamically-formed BH binaries with orbits similar to Gaia BH3 are likely to be discovered in Gaia DR4.
@@ 1 Introduction
Gaia Collaboration et al. (2024) recently reported discovery of an astrometric binary containing a $0.8\,M_{\odot}$ star on the lower giant branch and a $33\,M_{\odot}$ dark companion that is presumably a black hole (BH). The system, which they named Gaia BH3, was discovered with pre-release DR4 astrometry, and its orbital motion was confirmed with (nearly) independent radial velocity measurements from the Gaia RVS spectrometer (Cropper et al., 2018).
@@
Gaia BH3 is the widest BH binary discovered to date, with an orbital period of $\sim 4250$ days and a semimajor axis of $\sim 16.5$ au. The eccentricity is relatively high, $e\approx 0.73$, such that the star and BH pass within 4.5 au of each other at periastron. How the system formed is uncertain. Gaia Collaboration et al. (2024) note that a typical red supergiant would not fit within the current orbit at periastron. They thus consider formation of the system from an isolated binary unlikely, and invoke dynamical exchange as a possible alternative. Balbinot et al. (2024) report that Gaia BH3 is a member of the ED-2 stellar stream (Dodd et al., 2023), lending credence to the dynamical formation hypothesis. Because ED-2 stars do not display light element abundance variations characteristic of globular clusters, Balbinot et al. (2024) propose that ED-2 formed from a low-mass cluster ($M_{\rm cluster}\lesssim 4\times 10^{4}\,M_{\odot}$).
Here, I consider the possibility of formation from an isolated binary in more detail. While the association of Gaia BH3 with ED-2 appears robust, the progenitor cluster’s initial density and mass are uncertain, leaving open the possibility that Gaia BH3 is a primordial binary as opposed to one formed by dynamical exchange. I explore (a) the expected mass and maximum radius of the BH progenitor, and (b) the range of possible orbits that could have produced the current orbit in the presence of possible impulsive mass loss and kicks to the BH during its formation. I ultimately conclude that, while orbits similar to Gaia BH3’s can naturally be produced from isolated binaries if BHs are born with modest natal kicks, this channel is disfavored for Gaia BH3 in particular because such natal kicks would have ejected the binary from the ED-2 progenitor cluster.
@@
The remainder of this paper is organized as follows. In Section 2, I describe evolutionary models for low-metallicity massive stars representative of the BH progenitor. Section 3 uses Monte Carlo simulations of BH formation to assess the likely initial separation if Gaia BH3 formed from an isolated binary. Section 4 argues that BH companions appear to be strongly overrepresented at low metallicity, and that additional nearby BHs orbited by low-metallicity stars likely remain to be discovered. Future prospects and broader implications are briefly discussed in Section 5.
@@ 2 The BH progenitor
Figure 1: MESA evolutionary models for low-metallicity massive stars. The model shown in black is the same in all panels: a $35\,M_{\odot}$ star with metallicity $\rm[M/H]=-2.2$ and no rotation. Colored lines show models varying metallicity (upper left), initial mass (upper right), initial rotation rate (lower left), and overshooting or mixing length (lower right). Diamonds, circles, and stars mark core hydrogen exhaustion, the onset of carbon burning, and core collapse. Dashed diagonal lines show constant radius. Shaded gray region marks $R>800\,R_{\odot}$, beyond which the star would have been unlikely to fit inside the pre-SN orbit of Gaia BH3 (Section 3). Most of the models exceed this limit at their maximum radius and thus would have overflowed their Roche lobes. The only exceptions are rapidly-rotating models that undergo chemically homogeneous evolution (CHE; lower left) and models with reduced overshooting (lower right).
To estimate the maximum radius and expected mass loss history of the BH progenitor, I calculated a small suite of evolutionary models using MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). The calculations closely follow the setup described by Klencki et al. (2020), with MESA updated to version r23.05.1. The most important ingredients are summarized below, and I refer to Klencki et al. (2020, 2021) for more details.
@@
To verify that the problem setup is consistent with the one used by Klencki et al. (2020), I began by reproducing their $35\,M_{\odot}$ model at metallicity $Z=0.00017$ and comparing to their published tracks. The model reaches a maximum radius of $1331\,R_{\odot}$ shortly before core collapse. Their corresponding track has a maximum radius $1340\,R_{\odot}$ at the same point, and its evolution in the HR diagram is nearly indistinguishable from that of the model calculated here.
@@
The shaded regions in Figure 1 show $R>800\,R_{\odot}$. In Section 3, I argue that this is the largest progenitor that could plausibly have fit inside the orbit of Gaia BH3 before the BH’s formation. Most of the models exceed this radius shortly before core collapse. A notable exception is provided by the two rapidly-rotating models shown in the lower left panel of Figure 1, which go through chemically homogeneous evolution (CHE). At high rotation rates, large-scale meridional circulations can efficiently mix helium into stellar envelopes, preventing the buildup of a chemical gradient and core/envelope dichotomy that leads to expansion in nonrotating models (Maeder, 1987; Yoon & Langer, 2005; Woosley & Heger, 2006). Under CHE, stars evolve from the main sequence toward the helium main sequence, where they have effective temperatures above $10^{5}\,\rm K$ and radii smaller than $1\,R_{\odot}$. These models eventually ignite carbon burning and terminate their evolution without ever reaching $R>10\,R_{\odot}$. At the low metallicities considered here, CHE is predicted to occur with initial rotation rates of $\Omega/\Omega_{\rm crit}\geq 0.45$.
One of the main differences between the models shown here and higher-metallicity massive star models is that low-metallicity models remain compact during core helium burning: although they are red supergiants just before core collapse, they spend the large majority of their post-main sequence evolution as blue or yellow supergiants with $R\lesssim 100\,R_{\odot}$, only expanding to red supergiant dimensions during their final $\sim 1000$ years (see Klencki et al., 2020, for further discussion). This is illustrated in Figure 2, which shows the late-stage radius evolution for models with a range of masses (top panel) and metallicities (bottom panel). The fiducial model increases its radius by a factor of 2 in its final 800 years. This behavior is not found in higher metallicity models, even at “low” metallicites comparable to those found in the SMC and LMC. The upshot is that if Gaia BH3 formed as a primordial binary, there may not have been sufficient time for tides to circularize the orbit when the BH progenitor was a red supergiant.
Figure 2: Radius evolution of MESA models in the final years before core collapse. Mass is varied in the top panel, and metallicity in the bottom panel. At low metallicity, the models expand significantly in the final $\sim 1000$ yr of their evolution, only appearing as red supergiants shortly before core collapse. Expansion occurs earlier at higher metallicity. This means that wide low-metallicity binaries are less likely to be tidally circularized before the formation of a BH.
3 Constraints on the progenitor orbit
I use Monte Carlo simulations to study the combinations of pre-supernova [^1] Here “supernova” loosely refers to the death of the massive star and formation of the BH. Whether there was a successful explosion or a long-lived luminous transient is uncertain. (SN) orbits and natal kicks that could have produced an orbit similar to that of Gaia BH3. This approach follows Tauris et al. (2017) and El-Badry et al. (2024b); see Tauris & van den Heuvel (2023) for additional details about SN kicks. In brief, I simulate a large number ($N=10^{7}$) of orbits and kicks, assume that the SN occurs at a random phase with a kick oriented in a random direction, and predict post-SN parameters for all orbits. Finally, I analyze the initial parameters of the simulations for which the final period and eccentricity are close to the observed values for Gaia BH3: $P_{\rm orb,final}=4000-4500$ d, and $e_{\rm final}=0.68-0.78$. Narrowing these ranges reduces the number of surviving Monte Carlo samples but has no significant effect on their distribution.
Figure 3: Constraints on kicks and the pre-SN orbit of Gaia BH3-like binaries. I assume that the binary had a circular orbit before the SN and that the BH received a kick with velocity $v_{\rm kick}$ when it formed. I simulate $10^{7}$ pre-SN orbits with a uniform period distribution $P_{\rm orb}\sim\mathcal{U}(0,2P_{\rm orb,BH3})$, a pre-SN mass distribution $M_{2,\rm\,init}/M_{\odot}\sim\mathcal{U}(33,100)$, and a kick velocity distribution $v_{\rm kick}/({\rm km\,s^{-1}})\sim\mathcal{U}(0,100)$. Black histograms show the properies of all the systems whose post-SN orbits have periods and eccentricities similar to Gaia BH3 (upper left). Red histograms show the subset of these that result in a post-SN systemic velocity $v_{\rm sys}<10\,\rm km\,s^{-1}$, as required to remain bound to the ED-2 stream progenitor. Orbits with a broad range of pre-SN separations can produce the observed period and eccentricity. However, most of these require significant kicks and result in $v_{\rm sys}>10\,\rm km\,s^{-1}$. Simulations with $v_{\rm sys}<10\,\rm km\,s^{-1}$ can only produce the observed period and eccentricity for $a_{\rm init}\lesssim 1200\,R_{\odot}$. The BH progenitor would overflow its Roche lobe in such orbits for $R\gtrsim 800\,R_{\odot}$.
3.1 Circular orbits
I first consider pre-SN orbits that have zero eccentricity, as would be expected if tides circularized the orbit when the BH progenitor expanded. In this case, the system’s current eccentricity must be a result of natal kicks and/or instantaneous mass loss during the SN. I use the formalism from Brandt & Podsiadlowski (1995) to predict post-SN parameters.
I begin with a uniform distribution of pre-SN orbital periods between 0 and 8500 d (twice the currently observed orbital period of Gaia BH3). I simulate kicks assuming orientations distributed uniformly on the sky and velocities distributed uniformly between 0 and 100 $\rm km\,s^{-1}$. I take a uniform distribution of $\Delta m$, the mass loss of the BH progenitor during the SN, between 0 and 67 $M_{\odot}$, implying a progenitor mass of 33-100 $M_{\odot}$. This is not a population synthesis simulation – pre-SN orbits are unlikely to be uniformly distributed in all parameters – but should be viewed as an exploration of what combinations of orbits, kicks, and mass loss could have produced the observed orbit.
The results of this experiment are shown in Figure 3. Black histograms show all simulations that produce post-SN periods and eccentricities similar to Gaia BH3 today, while red histograms show the subset of those that predict a post-SN systemic velocity $v_{\rm sys}<10\,\rm km\,s^{-1}$. This is a conservative estimate of the largest post-SN systemic velocity for which Gaia BH3 could plausibly have remained bound to the low-mass cluster that formed the ED-2 stream.
When there is no restriction on $v_{\rm sys}$ (black histograms), initial separations ranging from $\sim 800$ to $\sim 6000\,R_{\odot}$ can produce the observed orbit. Initial separation closer than $\sim 800\,R_{\odot}$ cannot, because widening these to the separation observed today via a kick or mass loss would result in a higher final eccentricity than is observed. Wider initial separations could match the orbit for suitably chosen kicks, but these are excluded by the prior of $P_{\rm orb,\,init}<8500$ d.
The upper right panel of Figure 3 shows the pre-SN separations and post-SN systemic velocities of all simulations that match the observed period and eccentricity. The combination of low $v_{\rm sys}$ ($\lesssim 20\,\rm km\,s^{-1}$) and large initial separation ($\gtrsim 1200\,R_{\odot}$) is ruled out because such models produce too low eccentricities. The red histograms show that among simulations with $v_{\rm sys}<10\,\rm km\,s^{-1}$, only relatively tight pre-SN orbits with $a_{\rm init}<1200\,R_{\odot}$ can match the observed period and eccentricity. These calculations all have weak kicks ($v_{\rm kick}\lesssim 10\,\rm km\,s^{-1}$) and substantial mass loss, such that the final eccentricity is mainly a result of the Blaauw (1961) kick. Initial masses $M_{2,\rm init}\gtrsim 70\,M_{\odot}$ are ruled out because they become unbound due to mass loss.
The largest star that could fit inside a pre-SN orbit of semimajor axis $a_{\rm init}$ has radius $R_{\rm max}=a_{\rm init}f_{q}$, where $f_{q}(M_{1}/M_{2})$ is given by Eggleton (1983) and varies from 0.68 to 0.71 over $M_{1}=35-70\,M_{\odot}$, assuming $M_{2}=0.8\,M_{\odot}$. This means that for the pre-SN orbits that can produce the observed period and eccentricity of Gaia BH3 without escaping the ED-2 progenitor, the ±$\pm 1\sigma$ range of maximum radii is $R_{\rm max}\approx 680_{-87}^{+73}\,R_{\odot}$. As shown in Section 2, most plausible models for the progenitor of the BH reach maximum radii larger than this value, so it seems unlikely that the present orbit of Gaia BH3 could have resulted from isolated binary evolution if the pre-SN orbit was circular.
3.2 Eccentric orbits
Figure 4: Similar to Figure 3, but for eccentric pre-SN orbits. I simulate $10^{7}$ orbits with initial eccentricity $e_{\rm init}\sim\mathcal{U}(0,1)$ and the SN occurring at a random phase. Other parameters are distributed as in Figure 3. $d_{\rm peri,\,init}$ is the separation between the BH progenitor and companion star at periastron in the pre-SN orbit. Most orbits with $v_{\rm sys}<10\,\rm km\,s^{-1}$ (red) have $d_{\rm peri,\,init}\lesssim 1100\,R_{\odot}$ and thus would overflow their Roche lobes if the BH progenitor reached $R\gtrsim 800\,R_{\odot}$.
I next consider pre-SN orbits that are eccentric. These may be expected if the BH progenitor only expands late in its evolution, leaving insufficient time for tidal circularization (e.g. Figure 2). I again simulate $N=10^{7}$ orbits, now with a uniform distribution of eccentricities between 0 and 1. I assume that the SN occurs at a random phase, corresponding to a uniform distribution of the pre-SN mean anomaly, and use the formalism from Hurley et al. (2002) to predict the post-SN orbital eccentricity, period, and center-of-mass velocity.
The results are shown in Figure 4. In this case, I show the distributions of pre-SN periastron distance, $d_{\rm peri,\,init}$, rather than semimajor axis, since this is the key parameter for determining whether the BH progenitor would have overflowed its Roche lobe. As in the circular case, a broad range of initial orbits are possible for arbitrary kicks, but a relatively narrow range of $d_{\rm peri,\,init}$ can produce the observed period and eccentricity subject to the restriction of $v_{\rm sys}<10\,\rm km\,s^{-1}$. Most of these orbits have weak kicks and little mass loss, and thus are similar to the orbit of Gaia BH3 today, which has $d_{\rm peri}\approx 800\,R_{\odot}$. Orbits with significantly larger $d_{\rm peri,\,init}$ and weak kicks end up with longer periods and/or lower eccentricities than the values observed for Gaia BH3 today.
Among orbits that reproduce the observed eccentricity and period of Gaia BH3 with $v_{\rm sys}<10\,\rm km\,s^{-1}$, 90% have $d_{\rm peri,\,init}<1120\,R_{\odot}$, such that they would overflow their Roche lobes at periastron for radii $R>775\,R_{\odot}$. 99% would overflow their Roche lobes at periastron for $R>1125\,R_{\odot}$. Given that the MESA models shown in Figure 1 predict maximum radii of $\gtrsim 1000\,R_{\odot}$, I conclude that the BH progenitor is unlikely to have avoided Roche lobe overflow in any orbit that could have produced the current orbit of Gaia BH3, unless it formed through chemically homogeneous evolution or had a maximum radius significantly smaller than predicted by the evolutionary models shown in Figure 1.